# A Framework for Double-Blind Federated Adaptation of Foundation Models

### v1

@@ 4.2 Encrypted Inference
The overall encrypted inference protocol can be summarized as follows. At the beginning of the collaboration, each client $k$ encrypts (using its public key) each input in its local dataset based on a FHE encryption scheme and sends the encrypted inputs $\{\mathcal{E}(\mathbf{x}_{i})\}_{i=1}^{N_{k}}$ and the corresponding encrypted labels $\{\mathcal{E}(y_{i})\}_{i=1}^{N_{k}}$ to the server. The server applies the embedding function on the encrypted data to obtain the input to the first attention layer $\{\mathcal{E}(\mathbf{b}_{0}^{i})\}_{i=1}^{N_{k}}$. Subsequently, for each FL round, the server randomly selects a batch of $n$ samples from this set, say $\mathcal{E}(\mathbf{B}_{0})=[\mathcal{E}(\mathbf{b}_{0}^{1}),\mathcal{E}(\mathbf{b}_{0}^{2}),\cdots,\mathcal{E}(\mathbf{b}_{0}^{n})]$, and performs encrypted inference using the FHE-friendly first attention block ^$\widehat{\mathcal{B}}_{\widehat{\psi}_{1}}$. The output of this block is permuted using randomly selected $\Pi_{1}$ before being transmitted to the client. The client decrypts these representations (using its private key), re-encrypts again (using its public key), and returns it to the server. These re-encrypted values are used as the input to the next attention block and the same encrypted inference process is repeated for all the other blocks $\ell=2,\cdots,L$.
@@
When the client receives the output of the final transformer attention block $\mathcal{E}(\mathbf{B}_{L})$, the decrypted representations are passed through the classification head and the final predictions are again encrypted to get $\mathcal{E}(\mathbf{\hat{Y}})=[\mathcal{E}(\hat{y}^{1}),\mathcal{E}(\hat{y}^{2}),\cdots,\mathcal{E}(\hat{y}^{n})]$. These encrypted predictions are sent back to the server for per-sample loss computation in the encrypted domain. The server computes the average (or total) loss over the entire batch and sends this encrypted average loss to the client. The client decrypts this average batch loss and uses it for local learning. During this entire encrypted inference protocol, all the operations on the server are carried out in the encrypted domain. Since the server does not have access to the client’s private key, the server learns no information about the client’s local data. On the other hand, the client receives a batch of intermediate representations (in plaintext) after each attention block, but samples in each batch are randomly permuted based on $\mathbf{\Pi}_{\ell}$. Since the client cannot access $\mathbf{\Pi}_{\ell}$, it cannot use a model extraction attack to learn any useful information about the FM held by the server, as long as the batch size $n$ is sufficiently large. This ensures that the proposed encrypted inference protocol is double-blind, even though it is interactive in nature.
Figure 3: Parallel adapter illustration.
4.3 Local Learning
@@
Though various model adaptation strategies are available, the proposed framework requires an adaptation method that does not require backpropagation of gradients through the FM. This leaves us with only two possible choices: transfer learning (where only the classification head is learned) and parallel adapter (where both the classification head and a side adapter are learned). In this work, we adopt the low-rank parallel adapter method proposed in [36] (see Figure 3). This method requires access to intermediate representations after every transformer attention block, which is readily available (in plaintext) through the encrypted inference protocol described in Section 4.2.
@@
where $\mathbf{W}_{\ell}^{d}$ and $\mathbf{W}_{\ell}^{u}$ are the down and up-projection matrices and $\alpha_{\ell}$ is the scaling factor at block $\ell$. Since the adapter function in Eq. 6 is applied to each individual sample, the permutation of samples within a batch does not affect this computation. However, the adapter output in Eq. 5 depends on values from two consecutive blocks, which have undergone different permutations. Hence, it is necessary to ensure consistent permutation of the inputs. When operating a batch of samples, Eq. 5 can be reformulated as:
@@
where $\mathbf{H}_{0}=\mathbf{B}_{L}$. Note that the client receives only a permutation of intermediate representations, i.e., $(\mathbf{B}_{\ell}\cdot\Pi_{\ell})$ and not the original $\mathbf{B}_{\ell}$, $\forall~{}\ell\in[1,L]$. Hence, to facilitate the computations associated with the parallel adapter, the server also sends $(\Pi_{\ell-1}^{-1}\cdot\Pi_{\ell})$ for all $\ell\in[2,L+1]$ as well as $(\Pi_{L}^{-1}\cdot\Pi_{1})$ to the client. When the client receives $(\mathbf{B}_{L}\cdot\Pi_{L})$, it can compute $\mathbf{H}^{{}^{\prime}}_{0}=(\mathbf{B}_{L}\cdot\Pi_{L})\cdot(\Pi_{L}^{-1}\cdot\Pi_{1})=(\mathbf{B}_{L}\cdot\Pi_{1})=(\mathbf{H}_{0}\cdot\Pi_{1})$. This can be directly used in Eq. 7 along with $(\mathbf{B}_{1}\cdot\Pi_{1})$ to compute $\mathbf{H}^{{}^{\prime}}_{1}=(\mathbf{H}_{1}\cdot\Pi_{2})$. Following the same logic, it is possible to compute $\mathbf{H}^{{}^{\prime}}_{\ell}=(\mathbf{H}_{\ell}\cdot\Pi_{\ell+1}),~{}\forall~{}\ell\in[1,L]$. When the server receives the final encrypted predictions from the client, it can permute the encrypted labels of the batch using $\Pi_{L+1}$ before computing the per-sample losses and aggregating them. It must be emphasized that revealing $(\Pi_{\ell-1}^{-1}\cdot\Pi_{\ell})$ for all $\ell\in[2,L+1]$ as well as $(\Pi_{L}^{-1}\cdot\Pi_{1})$ to the client does not leak any information about $\Pi_{\ell}$ as shown in Proposition 1. Finally, the client locally updates the parallel adapter and classification head in the plaintext domain based on the average loss received from the server employing the same procedure as in [39].
@@
Non-uniqueness of solutions: To recover ${\bm{A}}$, ${\bm{B}}$, and ${\bm{C}}$ uniquely, we would need the products ${\bm{P}}$, ${\bm{Q}}$, and ${\bm{R}}$ to uniquely determine a single set of matrices $({\bm{A}},{\bm{B}},{\bm{C}})$. However, for any valid solution $({\bm{A}},{\bm{B}},{\bm{C}})$ that satisfies ${\bm{P}}={\bm{A}}^{-1}{\bm{B}}$, ${\bm{Q}}={\bm{B}}^{-1}{\bm{C}}$, and ${\bm{R}}={\bm{C}}^{-1}{\bm{A}}$, we can construct alternative solutions by applying a left-multiplication with any permutation matrix ${\bm{S}}\in\mathbb{P}_{n}$. Define alternative matrices ${\bm{A}}^{\prime}={\bm{S}}{\bm{A}}$, ${\bm{B}}^{\prime}={\bm{S}}{\bm{B}}$, and ${\bm{C}}^{\prime}={\bm{S}}{\bm{C}}$, then:

### v2

@@ 4.2.1 Encrypted Inference
The overall encrypted inference protocol can be summarized as follows. At the beginning of the collaboration, each client $\mathcal{P}_{k}$ encrypts (using its public key) its local inputs and labels $\{\mathcal{E}(\mathbf{x}_{i}),\mathcal{E}(y_{i})\}_{i=1}^{N_{k}}$ and sends them to the server. The server applies the embedding function on the encrypted data to obtain the input to the first attention layer $\{\mathcal{E}(\mathbf{b}_{0}^{i})\}_{i=1}^{N_{k}}$. Subsequently, for each FL round, the server randomly selects a batch of $n$ samples from this set, say $\mathcal{E}(\mathbf{B}_{0})=[\mathcal{E}(\mathbf{b}_{0}^{1}),\mathcal{E}(\mathbf{b}_{0}^{2}),\cdots,\mathcal{E}(\mathbf{b}_{0}^{n})]$, and sequentially performs encrypted inference on each FHE-friendly block $\widehat{\mathcal{B}}_{\widehat{\psi}_{\ell}}$. After block $\ell$, the client decrypts these representations (using its private key), re-encrypts them again (using its public key), and returns them to the next block.
@@
When the client receives the output of the final transformer attention block $\mathcal{E}(\mathbf{B}_{L})$, the decrypted representations are passed through the classification head and the final predictions are again encrypted to get $\mathcal{E}(\mathbf{\hat{Y}})=[\mathcal{E}(\hat{y}^{1}),\mathcal{E}(\hat{y}^{2}),\cdots,\mathcal{E}(\hat{y}^{n})]$. These encrypted predictions are sent back to the server for per-sample loss computation in the encrypted domain. The server computes the batch loss and sends this encrypted average loss to the client. The client decrypts this loss and uses it for local learning. Throughout, all the operations on the server are carried out in the encrypted domain. Since the server does not have access to the client’s private key, the server learns no information about the client’s local data. On the other hand, the client receives a batch of intermediate representations (in plaintext) after each attention block.
4.2.2 Local Learning
@@
Though various model adaptation strategies are available, the proposed framework requires an adaptation method that does not require backpropagation of gradients through the FM. This leaves us with only two possible choices: transfer learning (where only the classification head is learned) and parallel adapter (where both the classification head and a side adapter are learned). In this work, we adopt the low-rank parallel adapter method proposed in [41] (see Figure 4). This method requires access to intermediate representations after every transformer attention block, which is readily available (in plaintext) through the encrypted inference protocol described in Section 4.2.1.
Figure 4: Illustration of the parallel adapter design.
@@
where $\mathbf{W}_{\ell}^{d}$ and $\mathbf{W}_{\ell}^{u}$ are the down- and up-projection matrices and $\alpha_{\ell}$ is the scaling factor at block $\ell$. Finally, the client locally updates the parallel adapter and classification head in the plaintext domain based on the average loss received from the server, employing the same procedure as in [45].
4.2.3 Secure Aggregation
To ensure the secure aggregation of parameter updates from clients, we leverage secure multi-party computation (MPC) [6]. This approach enables the aggregation server to compute the average of client updates without gaining access to the individual updates themselves. In BlindFed, the local adapters and classifiers are securely aggregated via FedAvg to obtain the global adapter and classification head.
4.3 Model Privacy Boosting
The downsides of performing encrypted inference over one attention block at a time are two-fold. Firstly, it increases the communication cost because encrypted intermediate outputs are exchanged between the client and the server after every block in every FL round. Since communication efficiency is not one of our core constraints, we consider the increased communication cost as a limitation, not a red flag. However, since the intermediate representations $\mathbf{b}_{\ell}$ after every attention block are accessible to the client in plaintext form, a malicious client could use $(\mathbf{b}_{\ell-1},\mathbf{b}_{\ell})$ pairs for multiple training samples to mount a model extraction attack [34] and learn the parameters of each transformer block. This clearly violates the model privacy constraint. To circumvent this problem and preserve model privacy, we introduce two changes to the online adaptation stage, namely, sample-level permutation and stochastic block sampling.
4.3.1 Sample-level Permutation
Each communication round processes a batch of samples. Let $\mathcal{E}(\mathbf{B}_{\ell})=[\mathcal{E}(\mathbf{b}_{\ell}^{1}),\mathcal{E}(\mathbf{b}_{\ell}^{2}),\cdots,\mathcal{E}(\mathbf{b}_{\ell}^{n})]$ be a batch of encrypted intermediate representations corresponding to a client, where $n$ is the batch size. Before sending these representations to the client, the server applies a $n\times n$ permutation matrix $\mathbf{\Pi}_{\ell}$ and sends only the permuted batch $\mathcal{E}(\mathbf{B}_{\ell})\cdot\mathbf{\Pi}_{\ell}=[\mathcal{E}(\mathbf{b}_{\ell}^{\pi(1)}),\mathcal{E}(\mathbf{b}_{\ell}^{\pi(2)}),\cdots,\mathcal{E}(\mathbf{b}_{\ell}^{\pi(n)})]$ to the client. Here, $\pi(1),\pi(2),\cdots,\pi(n)]$ represents a random permutation of $1,2,\cdots,n]$. This permutation matrix $\mathbf{\Pi}_{\ell}$ can be randomly selected for each block $\ell$ in each communication round. Thus, the client never sees corresponding pairs $(\mathbf{b}_{\ell-1}^{i},\mathbf{b}_{\ell}^{i})$ for any training sample $i$ in the batch, ensuring some protection against model extraction attacks.
Because the adapter in Eq. 7 is applied per sample, the permutation of samples within a batch does not affect this computation. However, the adapter output in Eq. 6 depends on values from two consecutive blocks, which have undergone different permutations. Hence, it is necessary to ensure consistent permutation of the inputs. When operating on a batch of samples, Eq. 6 can be reformulated as:
@@ 4.3.1 Sample-level Permutation
where $\mathbf{H}_{0}=\mathbf{B}_{L}$. Note that the client receives only a permutation of intermediate representations, i.e., $(\mathbf{B}_{\ell}\cdot\Pi_{\ell})$ and not the original $\mathbf{B}_{\ell}$, $\forall~\ell\in[1,L]$. Hence, to facilitate the computations associated with the parallel adapter, the server also sends $(\Pi_{\ell-1}^{-1}\cdot\Pi_{\ell})$ for all $\ell\in[2,L]$ as well as $(\Pi_{L}^{-1}\cdot\Pi_{1})$ to the client. When the client receives $(\mathbf{B}_{L}\cdot\Pi_{L})$, it can compute $\mathbf{H}^{{}^{\prime}}_{0}=(\mathbf{B}_{L}\cdot\Pi_{L})\cdot(\Pi_{L}^{-1}\cdot\Pi_{1})=(\mathbf{B}_{L}\cdot\Pi_{1})=(\mathbf{H}_{0}\cdot\Pi_{1})$. This can be directly used in Eq. 8 along with $(\mathbf{B}_{1}\cdot\Pi_{1})$ to compute $\mathbf{H}^{{}^{\prime}}_{1}=(\mathbf{H}_{1}\cdot\Pi_{2})$. Following the same logic, it is possible to compute $\mathbf{H}^{{}^{\prime}}_{\ell}=(\mathbf{H}_{\ell}\cdot\Pi_{\ell+1}),~\forall~\ell\in[1,L]$. When the server receives the final encrypted predictions from the client, it can permute the encrypted labels of the batch using $\Pi_{L+1}$ before computing the per-sample losses and aggregating them. It must be emphasized that revealing $(\Pi_{\ell-1}^{-1}\cdot\Pi_{\ell})$ for all $\ell\in[2,L]$ as well as $(\Pi_{L}^{-1}\cdot\Pi_{1})$ to the client does not leak any information about $\Pi_{\ell}$ as shown in Proposition 1.
@@
See Appendix A.
4.3.2 Stochastic Block Sampling (SBS)
Sample-level permutation ensures that samples in each batch are randomly permuted based on $\mathbf{\Pi}_{\ell}$. Because the client never sees $\mathbf{\Pi}_{\ell}$, it is not straightforward to mount a model extraction attack when the batch size $n$ is sufficiently large. However, intermediate representations of the same sample by two successive transformer blocks are likely to have higher similarity than representations from different samples (Appendix B). This similarity can be exploited to recover individual permutation matrices or, at the very least, reduce the brute-force search space. To mitigate this risk, we introduce stochastic block sampling (SBS): at each server-side forward pass, we return only a subset of block outputs and set the rest to zero, so the full sequence of representations is never revealed.
Figure 5: Stochastic block sampling strategy.
A key consideration in this strategy is avoiding the sampling of consecutive (neighboring) blocks, as this could still enable similarity-based attacks. As shown in Figure 5 (and Appendix B), feature similarity is negligible when blocks are separated by at least one layer. We therefore use a structured sampling process: (i) if block $\ell$ is sampled (state $1$), the next block $\ell+1$ is not sampled (probability $1$); (ii) if block $\ell$ is not sampled (state 0), the next block $\ell+1$ is sampled with a probability of $0.5$. Thus, the proposed model privacy boosting techniques ensure that the encrypted inference protocol is double-blind, even though the intermediate representations are exposed in plaintext form.
Datasets Methods Is double-
blind?
Centralized Federated $(K=5)$
CIFAR-10 Full fine-tuning $0.9635$ $0.9759$ $0.9725$ $0.8857$
LoRA $0.9592$ $0.9736$ $0.9718$ $0.8979$
Adapter tuning $0.8992$ $0.8681$ $0.8539$ $0.6754$
Linear probing $0.9226$ $0.9203$ $0.9191$ $0.7447$
BlindFed $0.9428$ $0.9471$ $0.9413$ $\mathbf{0.8540}$
BlindFed + SBS $\mathbf{0.9443}$ $\mathbf{0.9486}$ $\mathbf{0.9427}$ $0.8489$
CIFAR-100 Full fine-tuning $0.8361$ $0.8684$ $0.8611$ $0.7882$
LoRA $0.8349$ $0.8593$ $0.8568$ $0.7647$
Adapter tuning $0.6594$ $0.6495$ $0.6396$ $0.4489$
Linear probing $0.7476$ $0.7486$ $0.7414$ $0.5317$
BlindFed $\mathbf{0.7930}$ $\mathbf{0.7929}$ $\mathbf{0.7808}$ $\mathbf{0.6620}$
BlindFed + SBS $0.7869$ $0.7861$ $0.7789$ $0.6584$
SVHN Full fine-tuning $0.9680$ $0.9763$ $0.9692$ $0.7601$
LoRA $0.9659$ $0.9656$ $0.9709$ $0.7545$
Adapter tuning $0.5201$ $0.5251$ $0.4785$ $0.3325$
Linear probing $0.5938$ $0.5879$ $0.5732$ $0.3385$
BlindFed $0.9232$ $\mathbf{0.9329}$ $0.9249$ $0.7431$
BlindFed + SBS $\mathbf{0.9257}$ $0.9298$ $\mathbf{0.9256}$ $\mathbf{0.7434}$
Table 1: Comparison of accuracy achieved by our proposed method against baseline approaches on three datasets (CIFAR-10, CIFAR-100, and SVHN) in both centralized and federated learning scenarios. Federated experiments involve five clients $(K=5)$ with data partitioned using a Dirichlet distribution at varying levels of heterogeneity $(\alpha=100,1,0.01)$. The best-performing results among double-blind algorithms are highlighted in bold.
5 Experiments and Results
5.1 Datasets
To validate and implement our approach, we utilize a well-known vision transformer (ViT) pretrained on ImageNet-1K (ViT-Base) as the FM. The public auxiliary datasets for distilling the FHE-friendly FM are: Tiny-Imagenet [31] is a subset of the larger ImageNet dataset, containing 200 different classes, each with 500 images, 100 validation/test images, totaling 120K images. Fed-ISIC2019 [57] is a multi-class dataset of dermoscopy images containing $23247$ images across $8$ melanoma classes, with high label imbalance. This dataset provides an opportunity to test our framework within the healthcare domain. For the distillation phase of our experiments, we exclusively use data from client 1.
The private downstream datasets are: CIFAR-10 and CIFAR-100 [30] datasets are standard benchmarks for image classification, containing $60000$ color images across $10$ and $100$ classes, respectively. SVHN [43] dataset is a benchmark for digit recognition, consisting of over $73257$ images of house numbers for training and $26032$ for testing. Fed-ISIC2019. The remaining data points for centers 2-6 are used in the fine-tuning experiments, aligning well with the federated setup, as the dataset is tailored for federated learning. For the centralized setup, all data points are pooled into a single client.
5.2 Experimental Setup
Public dataset Methods Is DB? Centralized Federated
Fed-ISIC2019
(center=0)
Full fine-tuning $0.7811$ $0.6752$
LoRA $0.7347$ $0.6844$
Adapter tuning $0.6601$ $0.5762$
Linear probing $0.6599$ $0.5856$
BlindFed $0.7090$ $0.6679$
BlindFed + SBS $\mathbf{0.7169}$ $\mathbf{0.6831}$
Tiny-Imagenet
Full fine-tuning $0.7817$ $0.6985$
LoRA $0.7330$ $0.6880$
Adapter tuning $0.6702$ $0.6074$
Linear probing $0.6372$ $0.5789$
BlindFed $0.7051$ $0.6481$
BlindFed + SBS $\mathbf{0.7127}$ $\mathbf{0.6581}$
Table 2: Performance comparison of our method with baseline approaches on the Fed-ISIC2019 dataset with five clients ($K=5$), using two auxiliary datasets: Fed-ISIC2019 (center=0) as an in-distribution (InD) dataset and Tiny-ImageNet as an out-of-distribution (OOD) dataset. DB refers to Double-Blind.
5.3 Results
Main results: In Table 1, the performance of our proposed method and baseline methods across three datasets (CIFAR-10, CIFAR-100, and SVHN) is evaluated under both centralized and federated learning settings. The federated learning experiment uses a Dirichlet partitioning strategy with varying levels of data heterogeneity, controlled by the Dirichlet concentration parameter $\alpha$ (ranging from $100$ to $0.01$). The results demonstrate that full fine-tuning achieves the highest accuracy across all datasets and settings, particularly excelling in more homogeneous federated scenarios, but it is computationally expensive and not double-blind. Linear probing maintains reasonable performance in homogeneous settings but fails drastically on SVHN and under extreme heterogeneity, confirming its limitations in adaptation. Our approach delivers robust and competitive performance, closely aligning with LoRA in accuracy while maintaining significantly lower computational demands (see Appendix D.2). Among the model privacy boosting techniques, sample-level permutation does not have any impact on the accuracy of the adapted model, but SBS may affect local learning because of missing intermediate representations. However, in practice, BlindFed+SBS demonstrates comparable performance to BlindFed without SBS, suggesting that SBS has minimal impact on adapted model performance while boosting model privacy. In some cases, these missing values add robustness to the learning process, leading to marginally better generalization performance.
Fed-ISIC2019 results: Table 2 compares the performance of our method and baselines on the Fed-ISIC2019 dataset with five centers. The auxiliary datasets used at the LSP include (1) Fed-ISIC2019 with only the first center (treated as an in-distribution dataset) and (2) Tiny-ImageNet (treated as an out-of-distribution dataset). The results demonstrate that knowledge transfer from the OOD dataset is effective for all the methods, highlighting that the auxiliary dataset used for offline distillation can be any available dataset.
Methods Is DB? $K=10$ $K=20$ $K=50$
Full fine-tuning $0.9739$ $0.9513$ $N/A$
LoRA $0.9661$ $0.9584$ $0.9482$
Adapter tuning $0.8696$ $0.8494$ $0.8165$
Linear probing $0.9167$ $0.9142$ $0.9007$
BlindFed $\mathbf{0.9446}$ $\mathbf{0.9422}$ $0.9287$
BlindFed + SBS $0.9425$ $0.9411$ $\mathbf{0.9388}$
Table 3: Scalability analysis of the proposed method to baseline approaches on the CIFAR-10 dataset, with varying number of clients $K\in\{10,20,50\}$ under a Dirichlet concentration parameter of $1.0$ for data partitioning. $N/A$ - one GPU is insufficient to run the experiment. DB refers to Double-Blind.
Scalability Results: Table 3 illustrates the scalability of our method and the baseline approaches on the CIFAR-10 dataset with an increasing number of clients $(K=\{10,20,50\})$ using Dirichlet partitioning $(\alpha=1.0)$ and a fixed batch size of 8. Full fine-tuning achieves the highest accuracy for $K=10$ and $K=20$ but becomes infeasible for $K=50$ due to GPU limitations (we use only one GPU for each of our experiments). Linear probing demonstrates stable performance, but our method outperforms linear probing in all the settings, balancing compute efficiency, scalability, and accuracy, and demonstrating its suitability for federated setups with a large number of clients.
Communication Overhead: In standard FL (FedAvg [40]), the communication cost depends on the foundation model (FM) size. In our work, ViT-Base is used as a FM consisting of ${\approx}86$M parameters requiring ${\approx}344$MB of bandwidth. In practice, LSPs often deploy larger FMs, e.g., ViT-Huge or even bigger models (${\approx}22$B parameters), which require ${\approx}88$GB of bandwidth. In contrast, BlindFed requires the transmission of an encrypted intermediate representation (IR) for each transformer block. In our work, an IR is a $577\times 768$ tensor, which requires $\mathbf{6.21MB}$ in plaintext and $C{=}\mathbf{17.33MB}$ after encryption (${\approx}\mathbf{2.8}\times$ expansion) (see Appendix D.5). Ignoring the tiny adapter update (0.250.25M params. ${\approx}1$MB), the total communication cost of BlindFed is $(N_{k}*L*C)$, transmitted in batches, where $N_{k}$ is the local training dataset size and $L$ is the no. of transformer blocks. For federated adaptation tasks, $N_{k}$ is expected to be small, and not all IRs need to be transmitted for SBS. Hence, the communication overhead of BlindFed will be significantly higher compared to FedAvg for smaller FMs but becomes comparable when adapting large FMs with limited local data (which is often the case in practice).
Other ablation studies and computational complexity analyses are reported in Appendix D.
6 Conclusions
This paper offers a promising framework for adapting foundation models for critical downstream applications without compromising on the data confidentiality and model privacy. However, the BlindFed framework is just a first step and needs to be improved in many dimensions before it is ready for practical adoption. Firstly, the high communication cost of BlindFed and high computational complexity at the server need to be mitigated. Secondly, more rigorous model privacy guarantees would be required before LSPs can expose valuable proprietary models to collaborators. Finally, the robustness of the framework in the presence of malicious collaborators should be carefully analyzed.
Acknowledgments
This material is partly based on work supported by the Office of Naval Research N00014-24-1-2168.
Supplementary Material
Appendix A Proof of Proposition 1
Proof of Proposition 1.
@@ Appendix A Proof of Proposition 1
Non-uniqueness of solutions:
To recover $\bm{A}$}, $\bm{B}$}, and $\bm{C}$} uniquely, we would need the products $\bm{P}$}, $\bm{Q}$}, and $\bm{R}$} to uniquely determine a single set of matrices $({\bm{A}},{\bm{B}},{\bm{C}})$. However, for any valid solution $({\bm{A}},{\bm{B}},{\bm{C}})$ that satisfies ${\bm{P}}={\bm{A}}^{-1}{\bm{B}}$, ${\bm{Q}}={\bm{B}}^{-1}{\bm{C}}$, and ${\bm{R}}={\bm{C}}^{-1}{\bm{A}}$, we can construct alternative solutions by applying a left multiplication with any permutation matrix ${\bm{S}}\in\mathbb{P}_{n}$. Define alternative matrices ${\bm{A}}^{\prime}={\bm{S}}{\bm{A}}$, ${\bm{B}}^{\prime}={\bm{S}}{\bm{B}}$, and ${\bm{C}}^{\prime}={\bm{S}}{\bm{C}}$, then: